# Gateformer: Advancing Multivariate Time Series Forecasting through Temporal and Variate-Wise Attention with Gated Representations

### v2

@@ Abstract
There has been a recent surge of interest in time series modeling using the Transformer architecture. However, forecasting multivariate time series with Transformer presents a unique challenge as it requires modeling both temporal (cross-time) and variate (cross-variate) dependencies. While Transformer-based models have gained popularity for their flexibility in capturing both sequential and cross-variate relationships, it is unclear how to best integrate these two sources of information in the context of the Transformer architecture while optimizing for both performance and efficiency. We re-purpose the Transformer architecture to effectively model both cross-time and cross-variate dependencies. Our approach begins by embedding each variate independently into a variate-wise representation that captures its cross-time dynamics, and then models cross-variate dependencies through attention mechanisms on these learned embeddings. Gating operations in both cross-time and cross-variate modeling phases regulate information flow, allowing the model to focus on the most relevant features for accurate predictions. Our method achieves state-of-the-art performance across 13 real-world datasets and can be seamlessly integrated into other Transformer-based and LLM-based forecasters, delivering performance improvements up to 20.7% over original models. Code is available at this repository: https://github.com/nyuolab/Gateformer.
@@ 1 Introduction
Time series forecasting is a ubiquitous problem with essential applications in various domains, such as traffic [Lv et al., 2014], energy [Zhu et al., 2023], weather [Angryk et al., 2020], healthcare [Kaushik et al., 2020], and finance [Chen et al., 2012]. In challenging multivariate time series, accurate predictions rely on effectively modeling both intra-series (cross-time) and inter-series (cross-variate) dependencies [Zhang and Yan, 2023]. Transformer architecture [Vaswani et al., 2017] revolutionized the modern natural language processing due to their ability to efficiently process sequential data and capture semantic relationships over long sequence lengths. Naturally, there has been a huge interest in transfering the capabilities of the attention mechanism to the time series domain [Wen et al., 2023].
Most Transformer-based forecasters [Kitaev et al., 2020, Li et al., 2021, Liu et al., 2022b] naively embed data points from all variables at the same time step into point tokens and model intra-series dependencies using attention mechanisms. However, these individual points lack explicit semantic information, often leading to ineffective temporal dependency modeling. As a result, simpler linear models can outperform these Transformer-based methods [Zeng et al., 2023]. PatchTST [Nie et al., 2023] addresses this limitation by expanding the receptive field, dividing time series into sequences of patches, and applying cross-time attention among patches to capture temporal dependencies, yielding promising results. However, these methods still overlook critical cross-variate relationships. To explicitly model cross-variate correlations, methods such as iTransformer [Liu et al., 2024b] have been proposed. However, iTransformer embeds each series into a coarse variate representation through linear projection, overlooking intra-series dependencies.
@@
In this work, we propose Gateformer, a Transformer-based model designed to explicitly capture both intra-series (cross-time) and inter-series (cross-variate) correlations. In the cross-time dependency modeling phase, we independently encode intra-series correlations of each variate into variate representations. These representations are then integrated with global temporal embeddings captured by an MLP through a gating mechanism. This gating operation regulates information flow, producing robust variate representations that serve as the foundation for subsequent cross-variate dependency modeling.
To capture cross-variate dependencies, we apply self-attention to these variate representations, generating variate-interacting embeddings. Recent work [Han et al., 2024] has shown that modeling cross-variate relationships enhances model capacity but requires larger datasets for effective correlation capture. Consequently, approaches focused on cross-variate modeling, like iTransformer, excel on high-dimensional datasets but show moderate performance on low-dimensional ones. To address this limitation, variate-interacting embeddings are integrated with outputs from the cross-time modeling phase (embeddings without variate interaction) through a gate operation to dynamically regulate the degree of cross-variate correlation, ensuring robust performance across datasets of varying sizes. Our main contributions in this work can be summarized as follows:
@@
  • We propose Gateformer, which utilizes temporal and variate-wise attention with gated representations to enhance multivariate time series forecasting.
@@
  • Our method achieves state-of-the-art results across 13 real-world benchmark datasets, ranking top-1 in 91 out of 130 settings and top-2 in 122 settings for varying prediction lengths and metrics.
@@
  • The proposed framework integrates seamlessly with Transformer-based and LLM-based models, delivering up to 20.7% performance improvement over the original models. Its flexibility enables the development of foundational multivariate models based on this approach.
2 Related work
Point-wise Representations with Cross-time Attention
Many prior works embed multiple variates at the same timestamp into point tokens and apply attention mechanisms to capture temporal dependencies among them. Examples include Autoformer [Wu et al., 2021], FEDformer [Zhou et al., 2022], and Pyraformer [Liu et al., 2022b], which focus on optimizing the quadratic complexity of self-attention. Autoformer detects sub-series similarities with $O(L\log L)$ complexity, FEDformer leverages frequency-domain sparsity with $O(L)$ complexity, and Pyraformer uses pyramidal attention for short- and long-term dependencies with $O(L)$ complexity. However, these models often fail in multivariate forecasting due to the lack of explicit semantic meaning in point tokens and the loss of cross-variate correlations when merging all variables into single temporal tokens.
Patch-wise Representations with Cross-time Attention PatchTST [Nie et al., 2023] addresses the lack of local semantic information through patching. It divides time series into patches to expand the receptive field and applies self-attention to capture cross-time dependencies among patches. PatchTST processes each variate independently, allowing unique attention patterns to be learned for each series. This approach has consistently improved performance on benchmarks and inspired recent large-scale time series models [Zhou et al., 2023, Das et al., 2024, Jin et al., 2024, Liu et al., 2024c]. However, it focuses primarily on cross-time dependencies while neglecting critical cross-variate interactions.
Variate-wise Representations with Cross-variate Attention A notable example is iTransformer [Liu et al., 2024b], which expands the receptive field as an extreme form of patching. iTransformer encodes each variate’s time series into coarse variate-wise representations through linear projection and explicitly models cross-variate correlations among these representations. However, it does not account for cross-time (intra-series) dependencies.
Patch-wise Representations with Cross-time and Cross-variate Attention Crossformer [Zhang and Yan, 2023], a representative model, segments each variate’s series into a sequence of patches and captures both cross-time (intra-series) and cross-variate (inter-series) dependencies among them. However, modeling multivariate correlations at such a granular level can introduce unnecessary noise, potentially resulting in suboptimal forecasting performance.
Unlike previous works, our approach captures both cross-time and cross-variate correlations through variate-wise representations. To enhance the quality of these representations, we integrate gating mechanisms to regulate information flow, allowing the model to focus on the most relevant features for accurate predictions. More details are provided in Section 3.
@@ 3.1 Problem definition
Given multivariate historical observations $\mathbf{X}\in\mathbb{R}^{N\times T}$ , time series forecasting aims to predict the next $F$ time steps, denoted as $\mathbf{\hat{Y}}\in\mathbb{R}^{N\times F}$ , where $N$ is the number of variates (or channels) and $T$ is the look-back window length. The objective is to minimize the mean squared error between $\hat{\mathbf{Y}}$ and the ground truth $\mathbf{Y}\in\mathbb{R}^{N\times F}$.
3.2 Model Overview
As shown in Figure 2, the proposed approach leverages the Transformer framework to explicitly capture cross-time and cross-variate dependencies, which are essential for multivariate forecasting.
Cross-time Dependency Modeling Phase Each variate’s time series is processed independently through two distinct pathways to generate variate-wise representations: (1) Temporal Dependency Embeddings, which capture cross-time correlations using patching and temporal-wise attention; and (2) Global Temporal Embeddings, which encode global temporal patterns using an MLP. These complementary embeddings are integrated through a gating operation to form the final variate embeddings.
Cross-variate Dependency Modeling Phase Variate-wise attention is applied to the variate embeddings to capture multivariate correlations, resulting in variate-interacting embeddings. However, modeling cross-variate dependencies requires more training data and can struggle with smaller datasets. To address this limitation, we integrate the variate-interacting embeddings with variate-independent embeddings (a copy of outputs from the cross-time modeling phase) to regulate the degree of cross-variate correlation, which is especially crucial in the presence of data scarcity.
Output Projection Layer The output from the cross-variate modeling phase is passed through a linear layer to generate the final predictions.
3.3 Model Structure
3.3.1 Cross-time Dependency Modeling
Accurately capturing intra-series dependencies is essential for creating high-quality series representations, which serve as the basis for subsequent cross-series dependency modeling. To achieve this, we independently encode each variate’s series as a variate-wise (series-wise) token, motivated by the following:
  • Each series is processed independently through the self-attention module, enabling it to learn unique attention patterns.
  • This approach minimizes the risk of noise in one series affecting others.
Temporal Dependency Each series $\mathbf{x}^{(i)}\in\mathbb{R}^{T}$ is normalized using reversible instance normalization (RevIN) to achieve zero mean and unit standard deviation, mitigating distribution shifts between training and testing data [Kim et al., 2022]. To preserve local semantics and increase the receptive field, $\mathbf{x}^{(i)}$ is divided into non-overlapping patches $\mathbf{x}^{(i)}_{P}\in\mathbb{R}^{P\times L_{p}}$, where $P=\lfloor\frac{T-L_{p}}{S}\rfloor+2$ denotes the number of patches, $L_{p}$ is the patch size, and $S$ is the stride (distance between the starting points of consecutive patches). These patches are mapped to $\hat{\mathbf{x}}^{(i)}_{P}\in\mathbb{R}^{P\times d_{m}}$ through a trainable linear projector to match the Transformer’s hidden dimension $d_{m}$.
To model cross-time (intra-series) dependency, we perform self-attention over $\hat{\mathbf{x}}^{(i)}_{P}$ to model correlation between patches. Queries, keys, and values ($\mathbf{Q}^{(i)}_{P},\mathbf{K}^{(i)}_{P},\mathbf{V}^{(i)}_{P}\in\mathbb{R}^{P\times d_{k}}$) are computed through linear mappings on $\hat{\mathbf{x}}^{(i)}_{P}$, where $d_{k}$ is the projected dimension. The self-attention output $\mathbf{O}^{(i)}_{cross-time}\in\mathbb{R}^{P\times d_{k}}$ is obtained as:
$\mathbf{O}^{(i)}_{cross-time}=\textsc{Attention}(\mathbf{Q}^{(i)}_{P},\mathbf{K}^{(i)}_{P},\mathbf{V}^{(i)}_{P})$
$=\textsc{Softmax}\left(\frac{\mathbf{Q}^{(i)}_{P}\mathbf{K}^{(i)\top}_{P}}{\sqrt{d_{k}}}\right)\mathbf{V}^{(i)}_{P}$
At this stage, the temporal dependency is now captured in $\mathbf{O}^{(i)}_{cross-time}$. We then reshape $\mathbf{O}^{(i)}_{cross-time}$ to $\mathbf{O^{\prime}}^{(i)}_{cross-time}\in\mathbb{R}^{d^{\prime}}$ ($d^{\prime}=P\times d_{k}$) and pass it through a feed-forward network to generate the temporal dependency embedding $\mathbf{v}^{(i)}_{T}\in\mathbb{R}^{d_{m}}$.
Global Temporal Pattern Although partitioning time series into patches and modeling dependencies among them effectively captures cross-time dependencies, as demonstrated by the success of PatchTST [Nie et al., 2023], it inevitably leads to information loss since patches only represent portions of the original series. To mitigate this, each series $\mathbf{x}^{(i)}$ is passed through a shared feed-forward network to extract global temporal patterns shared across variates and generate the global temporal embeddings. Specifically, the network linearly transforms $\mathbf{x}^{(i)}$ into $\mathbf{v}^{(i)}_{G}\in\mathbb{R}^{d_{m}}$. Recent studies [Zeng et al., 2023, Das et al., 2023, Li et al., 2023] have shown that MLPs can effectively capture temporal patterns often shared across different time series.
3.3.2 Gated Fusion
To improve the quality of variate embeddings, a gating mechanism is used to integrate representations encoding temporal dependencies $\mathbf{v}^{(i)}_{T}$ with those capturing global temporal patterns $\mathbf{v}^{(i)}_{G}$. This allows the model to focus on relevant information for accurate predictions. Formally, both embeddings are projected into a latent space, combined element-wise, and passed through a sigmoid function to compute the gate value. The embeddings are then integrated as follows:
$Gate=\sigma(\mathbf{v}^{(i)}_{T}\boldsymbol{W}_{g1}+\mathbf{v}^{(i)}_{G}\boldsymbol{W}_{g2})$ (1)
$\mathbf{s}^{(i)}=Gate\odot\mathbf{v}^{(i)}_{T}+(1-Gate)\odot\mathbf{v}^{(i)}_{G}$ (2)
where $\boldsymbol{W}_{g1}$ and $\boldsymbol{W}_{g2}$ are learnable parameters, $\sigma(\cdot)$ is the sigmoid function, and $\mathbf{s}^{(i)}\in\mathbb{R}^{d_{m}}$ is the final series representation. This mechanism produces a robust variate representation that encodes both temporal dependencies and global temporal patterns, providing a strong basis for cross-variate (cross-series) dependency modeling.
3.3.3 Cross-variates Dependency Modeling
A variate-wise attention module is used to model cross-variate dependencies. Given the variate embeddings $\mathbf{S}=\{\mathbf{s}^{(0)},\ldots,\mathbf{s}^{(N)}\}\in\mathbb{R}^{N\times d_{m}}$, the self-attention mechanism projects $\mathbf{S}$ into queries, keys, and values $\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{N\times d_{m}}$. Following the standard self-attention procedure, the output $\mathbf{O}_{cross-variate}\in\mathbb{R}^{N\times d_{m}}$ is computed as:
$\mathbf{O}_{cross-variate}=\textsc{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})$
$=\textsc{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}$
While the multivariate correlations are encoded in $\mathbf{O}_{cross-variate}$, this approach requires more training data to effectively capture cross-variate dependencies, leading to moderate performance on smaller datasets. To mitigate this, a gating mechanism is used again to integrate the embeddings from channel-interacting $\mathbf{O}_{cross-variate}$ and channel-independent perspectives $\mathbf{S}$, ensuring robust forecasting results across both large and small datasets.
3.3.4 Output Projection Layer
Directly producing long-term forecasts using a linear layer has proven highly effective [Nie et al., 2023, Zeng et al., 2023]. Thus, a linear head is applied to $\mathbf{O}_{cross-variate}\in\mathbb{R}^{N\times d_{m}}$ to generate the prediction $\mathbf{\hat{Y}}\in\mathbb{R}^{N\times F}$. The mean and standard deviation used to normalize the input series are restored to the output predictions. Mean squared error (MSE) is used to measure the discrepancy between predictions $\mathbf{\hat{Y}}$ and ground truth $\mathbf{Y}$.
Table 1: Statistics of the thirteen datasets used in our experiments
Datasets Variates Timesteps Frequency
ETTh1&2 7 17,420 hourly
ETTm1&2 7 69,680 15 min
Traffic 862 17,544 hourly
Electricity 321 26,304 hourly
Weather 21 52,696 10 min
Exchange 8 7,588 daily
Solar-Energy 137 52,560 10 min
PEMS03 137 26,208 5 min
PEMS04 358 16,992 5 min
PEMS07 307 28,224 5 min
PEMS08 883 17,856 5 min
@@ 3.3.4 Output Projection Layer
Table 2: Multivariate forecasting results with prediction lengths $T\in\{96,192,336,720\}$ and fixed look-back length $L=96$. Results are averaged from all prediction lengths. A lower value indicates better performance. Full results are listed in the Appendix G.
@@ 4 EXPERIMENTS
We extensively evaluate the proposed model on both long-term and short-term time series forecasting tasks, comparing its performance against nine state-of-the-art (SOTA) methods. An ablation study highlights the contribution of each model component. Additionally, our experiments demonstrate that the proposed method enhances both Transformer-based and LLM-based forecasters, delivering substantial performance improvements over the original models.
Dataset We use thirteen real-world datasets in the experiments, including Traffic, Electricity, Weather, Exchange, ETT (4 subsets) from Autoformer [Wu et al., 2021], Solar-Energy from LSTNet [Lai et al., 2018], and PEMS (4 subsets) from SCINet [Liu et al., 2022a]. For a fair comparison, we follow the standard protocol [Wu et al., 2023, Liu et al., 2024b], splitting the datasets into training, validation, and test sets using ratios of 6:2:2 for ETT and PEMS datasets, and 7:1:2 for the others. The splits are strictly chronological to ensure no data leakage occurs. The characteristics of these datasets are shown in Table 1, with additional details provided in Appendix A.1.
Table 4: Performance improvements achieved by our framework. The results are averaged across four prediction lengths $T\in\{96,192,336,720\}$ with a look-back window of $L=96$. Relative reductions in MSE and MAE (Promotion) are reported. Full results are provided in the Appendix D.
Models Autoformer Flowformer GPT4TS
MSE MAE MSE MAE MSE MAE
ECL Original 0.227 0.338 0.267 0.359 0.217 0.308
+ Our framework 0.198 0.283 0.171 0.263 0.177 0.267
Promotion 12.8% 16.3% 36.0% 26.7% 18.4% 13.3%
Weather Original 0.338 0.382 0.286 0.308 0.279 0.297
+ Our framework 0.257 0.281 0.255 0.279 0.261 0.282
Promotion 24.0% 26.4% 10.8% 9.4% 6.4% 5.0%
Models Gateformer Models Trained from Scratch on Target Dataset
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE

### v3

# Gateformer: Advancing Multivariate Time Series Forecasting via Temporal and Variate-Wise Attention with Gated Representations

@@ Abstract
Transformer-based models have recently shown promise in time series forecasting, yet effectively modeling multivariate time series remains challenging due to the need to capture both temporal (cross-time) and variate (cross-variate) dependencies. While prior methods attempt to address both, it remains unclear how to optimally integrate these dependencies within the Transformer architecture for both accuracy and efficiency. We re-purpose the Transformer to explicitly model these two types of dependencies: first embedding each variate independently to capture temporal dynamics, then applying attention over these embeddings to model cross-variate relationships. Gating mechanisms in both stages regulate information flow, enabling the model to focus on relevant features. Our approach achieves state-of-the-art performance on 13 real-world datasets and can be integrated into Transformer-, LLM-based, and foundation time series models, improving performance by up to 20.7%. Code is available at this repository: https://github.com/nyuolab/Gateformer.
Machine Learning, ICML
@@ 1 Introduction
Time series forecasting plays a critical role in domains such as traffic (Lv et al., 2014), energy (Zhu et al., 2023), weather (Angryk et al., 2020), healthcare (Kaushik et al., 2020), and finance (Chen et al., 2012). In multivariate settings, accurate forecasting depends on modeling both intra-series (temporal) and inter-series (cross-variable) dependencies (Zhang & Yan, 2023). Transformers (Vaswani et al., 2017), originally successful in NLP, have been adapted for time series due to their ability to model long-range dependencies (Wen et al., 2023).
Many Transformer-based models (Kitaev et al., 2020; Li et al., 2021; Liu et al., 2022b) embed all variables at each time step into a single token and apply attention across time. However, these tokens often lack semantic depth, weakening temporal modeling—sometimes allowing simple linear models to outperform them (Zeng et al., 2023). PatchTST (Nie et al., 2023) addresses this by using patch-level temporal attention but overlooks cross-variable dependencies. iTransformer (Liu et al., 2024b) models cross-variable correlations via coarse embeddings, sacrificing fine-grained temporal detail.
@@
We propose Gateformer, a Transformer-based model that explicitly captures both temporal and cross-variate dependencies. Each variate is first encoded independently to capture temporal patterns, then integrated with global temporal context through a gating mechanism that regulates information flow. These representations serve as inputs for cross-variate modeling.
To model cross-variate dependencies, we apply self-attention over variate representations, producing variate-interacting embeddings. While such modeling improves capacity (Han et al., 2024), it may degrade performance on low-dimensional datasets. To address this, we fuse variate-interacting and non-interacting embeddings through a second gating mechanism, dynamically controlling cross-variate influence and ensuring robust performance across dataset scales. Our main contributions in this work can be summarized as follows:
@@
  • We introduce Gateformer, which combines temporal and variate-wise attention with gated representations to improve multivariate time series forecasting.
@@
  • Gateformer achieves state-of-the-art results on 13 real-world benchmarks, ranking top-1 in 91 and top-2 in 122 out of 130 settings.
@@
  • Our propsed framework seamlessly integrates with Transformer-based, LLM-based, and foundation time series models, achieving performance improvements of up to 20.7% and facilitating the development of foundational multivariate models.
@@ 2.1 Problem Definition
Given multivariate time series data $\mathbf{X}\in\mathbb{R}^{N\times T}$, the goal is to forecast the next $F$ steps, $\mathbf{\hat{Y}}\in\mathbb{R}^{N\times F}$, where $N$ is the number of variates and $T$ is the look-back window. The objective is to minimize the MSE between $\hat{\mathbf{Y}}$ and the ground truth $\mathbf{Y}$.
@@
Table 1: Multivariate forecasting results with prediction lengths $T\in\{96,192,336,720\}$ and fixed look-back length $L=96$. Results are averaged from all prediction lengths. A lower value indicates better performance. Full results are listed in the Appendix J.
@@
2.2 Model Overview
As shown in Figure 1, our model captures both temporal and cross-variate dependencies.
Cross-time Dependency Modeling. Each variate is processed independently to capture intra-series patterns. Two parallel paths are used: (1) patch-based temporal attention to capture local dependencies, and (2) a shared MLP to learn global temporal patterns. These are fused via a gating mechanism to form variate embeddings.
Cross-variate Dependency Modeling. Variate embeddings are passed through a variate-wise attention module to model inter-series relationships. To maintain performance on small datasets, these outputs are gated with original variate embeddings to dynamically control cross-variate influence.
Output Projection. Final embeddings are passed through a linear layer to produce predictions $\hat{\mathbf{Y}}$, followed by de-normalization. MSE is used for training.
2.3 Model Details
2.3.1 Cross-time Dependency Modeling
Each series $\mathbf{x}^{(i)}\in\mathbb{R}^{T}$ is normalized with RevIN (Kim et al., 2022), then divided into $P$ non-overlapping patches. Each patch is projected to a $d_{m}$-dimensional space and passed through self-attention:
$\mathbf{O}^{(i)}_{cross-time}=\textsc{Softmax}\left(\frac{\mathbf{Q}^{(i)}\mathbf{K}^{(i)\top}}{\sqrt{d_{k}}}\right)\mathbf{V}^{(i)}$
The attention output is reshaped and passed through an FFN to obtain temporal embeddings $\mathbf{v}^{(i)}_{T}\in\mathbb{R}^{d_{m}}$.
Global Temporal Pattern. To address the loss of global context from patching, each raw series is also processed by a shared MLP to obtain a global embedding $\mathbf{v}^{(i)}_{G}\in\mathbb{R}^{d_{m}}$.
2.3.2 Gated Fusion
We fuse $\mathbf{v}^{(i)}_{T}$ and $\mathbf{v}^{(i)}_{G}$ using a gate:
$Gate=\sigma(\mathbf{v}^{(i)}_{T}\boldsymbol{W}_{g1}+\mathbf{v}^{(i)}_{G}\boldsymbol{W}_{g2})$
$\mathbf{s}^{(i)}=Gate\odot\mathbf{v}^{(i)}_{T}+(1-Gate)\odot\mathbf{v}^{(i)}_{G}$
$\mathbf{s}^{(i)}$ serves as the final variate embedding.
2.3.3 Cross-variate Dependency Modeling
Given all variate embeddings $\mathbf{S}\in\mathbb{R}^{N\times d_{m}}$, we apply self-attention:
$\mathbf{O}_{cross-variate}=\textsc{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{m}}}\right)\mathbf{V}$
To ensure robustness across dataset sizes, the output is gated with the original $\mathbf{S}$ embeddings.
2.3.4 Output Projection Layer
We apply a linear head on $\mathbf{O}_{cross-variate}$ to obtain final predictions $\hat{\mathbf{Y}}\in\mathbb{R}^{N\times F}$. Predictions are de-normalized using stored mean and standard deviation, and MSE is used as the training loss.
Models
Autoformer
Flowformer
GPT4TS
Moment
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Electricity Original 96 0.201 0.317 0.215 0.320 0.197 0.290 0.204 0.296
192 0.222 0.334 0.259 0.355 0.201 0.292 0.207 0.299
336 0.231 0.338 0.296 0.383 0.217 0.309 0.219 0.310
720 0.254 0.361 0.296 0.380 0.253 0.339 0.256 0.341
Avg
0.227
0.338
0.267
0.359
0.217
0.308
0.221
0.311
+ Our framework 96 0.173 0.259 0.144 0.237 0.150 0.241 0.163 0.254
192 0.180 0.267 0.160 0.251 0.163 0.254 0.177 0.269
336 0.198 0.285 0.175 0.269 0.180 0.272 0.195 0.283
720 0.241 0.320 0.205 0.296 0.214 0.301 0.226 0.311
Avg
0.198
0.283
0.187
0.274
0.177
0.267
0.190
0.279
Weather Original 96 0.266 0.336 0.182 0.233 0.203 0.244 0.192 0.234
192 0.307 0.367 0.250 0.288 0.247 0.277 0.246 0.278
336 0.359 0.395 0.309 0.329 0.297 0.311 0.287 0.305
720 0.419 0.428 0.404 0.385 0.368 0.356 0.360 0.350
Avg
0.338
0.382
0.286
0.308
0.279
0.297
0.271
0.292
+ Our framework 96 0.175 0.218 0.171 0.213 0.175 0.214 0.184 0.223
192 0.219 0.258 0.220 0.258 0.228 0.260 0.225 0.259
336 0.281 0.300 0.275 0.297 0.282 0.301 0.281 0.301
720 0.353 0.350 0.353 0.349 0.359 0.352 0.356 0.349
Avg
0.257
0.281
0.255
0.279
0.261
0.282
0.261
0.283
@@ 3 Experiments
We evaluate our model on short- and long-term forecasting tasks against nine SOTA methods. Results show strong performance and notable gains when applied to Transformer- and LLM-based forecasters. Ablation studies highlight the contribution of each component.
Datasets. We use 13 real-world datasets, including Traffic, Electricity, Weather, Exchange, four ETT subsets (Wu et al., 2021), Solar-Energy (Lai et al., 2018), and four PEMS subsets (Liu et al., 2022a). Dataset details are provided in Appendix B.1.